OpenStudy (anonymous):

How do I do: (square root symbol) 6x^8y^9/ (square root symbol) 5x^2y^4

OpenStudy (zale101):

Since there is not keys for square root on your keyboard, then just type sqrt. Many people i see use sqrt as the symbol of square root, including me.

jhonyy9 (jhonyy9):

push on ,,EQUATION" with blue base there will get a table where you can selrcting the squre root symbol

jhonyy9 (jhonyy9):

\[\sqrt{x}\]

OpenStudy (anonymous):

Lol thanks for telling me how to add the square root symbol lol it'll help a lot but I need to know how to do the actusal equation so just pretend the symbol is there

OpenStudy (anonymous):

Actual*

jhonyy9 (jhonyy9):

\[\sqrt{6x^8y^9}/\sqrt{5x^2y^4}\]

jhonyy9 (jhonyy9):

hope that will help you

OpenStudy (anonymous):

I need to know how to SOLVE the equation lol

jhonyy9 (jhonyy9):

sqrt(6x^8y^9 6x^8y^9 6 6 ------------- = sqrt( ---------) = sqrt( -- * x^(8-2)*y^(9-4) ) = sqrt( -- * x^6*y^5 ) sqrt5x^2y^4 5x^2y^4 5 5 6 6 = sqrt( -- * (x^3)^2 *(y^2)^2 *y ) = x^3*y^2 *sqrt(--*y) 5 5 hope so much that you understand it right easy good luck bye